William is 2 times as old as Stephanie. 28 years ago, William was 9 times as old asStephanie.

Answer by macston(5194)

S=stephanie's age; W=William's age=2S
W-28=9(S-28) Substitute for W
2S-28=9S-252 Add 252 to each side.
2S+224=9S Subtract 2S from each side.
224=7S Divide each side by 7
32 years=S ANSWER 1: Stephanie is 32 years old.
W=2S=2(32years)=64 years ANSWER 2: William is 64 years old.